given the demand function Pd=25-Q2and supply function Ps = 2Q+1. Assuming pure competition, find the consumer and producer surplus.


1
Expert's answer
2021-04-21T07:57:11-0400

At equilibrium, Pd = Ps.

"25\u2212Q^{2}= 2Q+1."

On solving, we get Q* = 4, and P* = 9.

Consumer surplus is calculated as

"\u222b^{q}_0d(Q) dQ \u2212 P*Q*"

Putting values

"CS= \u222b^{4}_025\u2212Q^{2} dQ \u2212 (9)(4)"

"= [25Q − \frac{Q^{3}}{3}] − 36."

Since Q = 4

"= [100 \u2212 21.33] \u2212 36"

=$ "42.67."

Now, producer surplus is calculated as

"P*Q* \u2212\u222b^{q*}_0s(Q) dQ."

Putting values

"PS = (9)(4) \u2212 \u222b^{4}_02Q+1 dQ"

"= 36 − [\frac{2Q^{2}}{2}+ Q]."

Since Q = 4,

"PS = 36 \u2212 16 \u2212 4"

=$"16"